A wire of length 3.00m and radius 2.0mm is stretched by 30.0mm on application of a force of 1500Newton. Calculate the Young's modulus for the wire
1
Expert's answer
2020-09-21T06:22:04-0400

"\sigma=E\cdot\frac{\Delta l}{l} \to E=\frac{l}{\Delta l}\cdot\frac{F}{A}="


"=\frac{l}{\Delta l}\cdot\frac{F}{\pi r^2}=\frac{3}{0.03}\cdot\frac{1500}{3.14\cdot 0.002^2}=11.9(GPa)"